OpenStudy (anonymous):

Please Help :) Solve. x2 + 8x + 25 = 0

OpenStudy (anonymous):

x = -2.5

OpenStudy (anonymous):

or -25/10

OpenStudy (anonymous):

x^2 + 8x + 25 = 0 2x or x squared? for the first term.

OpenStudy (anonymous):

a. -4 ± i b. -2 ± 4i c. -4 ± 3i d. -3 ± 4i

OpenStudy (anonymous):

oh okay, quadratic equation.

OpenStudy (anonymous):

\[\frac{ -8 \pm \sqrt{-36} }{ 2 }\] \[\frac{ -b \pm \sqrt{b ^{2} - 4ac}}{ 2a }\] \[\frac{ -8 \pm \sqrt{64 - 4(1)(25)} }{ 2 }\] -8 +\- 6i all over 2 -4 +\- 3i answer is C

OpenStudy (anonymous):

Thank you @victorbojica :)

OpenStudy (anonymous):

Sorry it took so long, tried to give you a visual with the equations.

OpenStudy (anonymous):

thats okay :)